A 20. L container is filled with helium and the pressure is 150 atm and the temperature is 30.° C. How many 5.0 L balloons can be filled when the temperature is 22° C and the atmospheric pressure is 755 mmHg? 


1
Expert's answer
2021-01-27T04:39:27-0500

pV=nRT. n=pV/RT = 150 x 20/0.082 x 303 = 120,743782 moles.

n= 755 x 5000 / 295 x 62400 = 0,210506302 moles.

120.743782/0,210506302= 573